OpenStudy (anonymous):

What is the solution to the equation x – 9 = –4? a.-13 b.-36 c.-5 d.5

OpenStudy (anonymous):

how did u get that answer?

OpenStudy (anonymous):

-5-9=4

undeadknight26 (undeadknight26):

@Obvial plug n play: it is not -5.

OpenStudy (anonymous):

then it is what?

undeadknight26 (undeadknight26):

try to plug in 5 for x

undeadknight26 (undeadknight26):

since youa re taking 9 away from 5 u get -4.

OpenStudy (anonymous):

to postives to give a negative love

OpenStudy (anonymous):

two*

OpenStudy (anonymous):

add \(9\) to both sides

OpenStudy (anonymous):

dont give *

undeadknight26 (undeadknight26):

you can also do it that way which is actually smarter...lol @satellite73

OpenStudy (anonymous):

Sorry :( its D: 5

OpenStudy (anonymous):

\[x-9=-4\\ x=-4+9\\ x=-5\]

OpenStudy (anonymous):

oh okay

undeadknight26 (undeadknight26):

x = 5 you mean @satellite73 ?

OpenStudy (anonymous):

You have to add the nine to both sides to get x all by itself.

undeadknight26 (undeadknight26):

exactly :D

OpenStudy (anonymous):

That was my first answer to a question :( sorry still trying to figure out how to work this thing lol

undeadknight26 (undeadknight26):

its fun trust me @Monster143 Good luck :D

OpenStudy (anonymous):

haha thanks :)